# Charting the Landscape of Nefarious Uses of Generative Artificial Intelligence for Online Election Interference

### v3

@@ Introduction
The constant exposure to AI-generated misinformation and deepfakes can have a psychological toll on the populace. The bombardment of conflicting information can lead to information overload, where individuals struggle to process and make sense of the vast amounts of data they encounter daily: This situation can cause increased stress, anxiety, and a sense of helplessness among individuals (Fu et al., 2020; Matthes et al., 2020; Brooks et al., 2024). Additionally, the fear of being deceived by AI-generated content can lead to a general mistrust of digital information. This mistrust can hinder the adoption of beneficial technologies and stifle innovation, as people become wary of engaging with digital platforms (Ferrara, 2024a).
Table 3. Examples of Research on Nefarious Applications of AI for Election Interference
Study Main Findings
Ratkiewicz et al. (2011) Introduces the "Truthy" project, which detects and tracks political smears, astroturfing, misinformation, and other political abuse in social media.
Bessi and Ferrara (2016) Found that social bots – later associated with Russia’s IRA (Broniatowski et al., 2018; Badawy et al., 2019a) – were used to significantly influence the 2016 U.S. Presidential election discussions on Twitter.
Shao et al. (2018) Identified that social bots amplify low-credibility content and contributed to the spread of misinformation during the 2016 U.S. election.
Im et al. (2020) Demonstrated that Russian troll accounts continue to operate and adapt their strategies post-2016 U.S. elections.
Haider et al. (2023) Developed a framework to demonstrate an approach to detect online political abuse in low-resource languages.
Stella et al. (2018) Showed that bots disproportionately amplify negative and inflammatory content, leading to greater exposure and potential influence on public opinion.
Chang et al. (2021) Explored how digital civic participation and misinformation campaigns affected the 2020 Taiwanese Presidential election.
Luceri et al. (2019b) Analyzed the behavior of bots and humans on social media during elections, finding that bots become more sophisticated and difficult to detect over time.
@@
Effective regulation and oversight are crucial in governing the use of GenAI, particularly in sensitive areas like elections. Governments and international bodies need to establish clear guidelines and standards for the ethical use of AI technologies. These regulations should include provisions for transparency, accountability, and the prevention of misuse (Floridi, 2019).
One approach is the implementation of AI ethics guidelines, such as the European Union’s Ethics Guidelines for Trustworthy AI, which emphasize principles like human agency, technical robustness, privacy, and accountability (Independent High-Level Expert Group on Artificial Intelligence, 2019). Additionally, regulatory frameworks should mandate the disclosure of AI-generated content to ensure that the public can identify and scrutinize such content appropriately (OpenAI, 2019).
@@
Advances in technology can provide robust solutions to detect and mitigate the impact of AI-generated misinformation. For instance, digital watermarking and forensic techniques can be employed to identify deepfakes and other synthetic media (Perov et al., 2020). These technologies can help verify the authenticity of digital content, making it easier to distinguish genuine from manipulated media.
Moreover, AI-driven detection systems can be developed to identify and neutralize botnets and fake accounts on social media platforms. These systems can use machine learning algorithms to detect patterns indicative of automated behavior, enabling platform operators to take swift action against malicious bots (Yang et al., 2022). Implementing such technologies requires collaboration between tech companies, researchers, and policymakers to ensure their effectiveness and scalability.
@@
Educating the public about the potential for AI-driven misinformation and how to critically evaluate information sources is crucial in combating the spread of false information. Public awareness campaigns can help individuals develop the skills needed to identify deepfakes, recognize signs of bot activity, and verify the credibility of news sources (Rudin, 2019).
Educational initiatives should be integrated into school curricula and adult education programs to promote media literacy. These programs can teach critical thinking skills and provide tools for assessing the reliability of information online (Ferrara, 2023b). Additionally, public service announcements and partnerships with media organizations can amplify these efforts, reaching a broader audience.
@@
Addressing the challenges posed by GenAI requires collaborative efforts among policymakers, technologists, and civil society. Multi-stakeholder initiatives can foster the development of best practices and shared standards for the ethical use of AI. For example, the Partnership on AI brings together diverse organizations to advance public understanding of AI and promote the responsible use of AI technologies (Heer, 2018).
Moreover, international cooperation is essential to address the global nature of online election interference. Governments and international organizations should work together to share information, coordinate responses, and develop joint strategies to counteract AI-driven threats. This collaboration can help build a unified front against the misuse of GenAI, ensuring that efforts are aligned and resources are effectively utilized (Taddeo and Floridi, 2018).
@@
Promoting ethical AI development involves ensuring that AI systems are designed with fairness, accountability, and transparency in mind. Developers and researchers should adhere to ethical guidelines and consider the societal impact of their work. This includes conducting thorough risk assessments, implementing bias mitigation strategies, and fostering inclusive design practices (Weidinger et al., 2022).
Ethical AI development also involves creating AI systems that are interpretable and explainable. Interpretable models allow users to understand how decisions are made, increasing trust in AI systems and enabling more effective oversight (Rudin, 2019). By prioritizing ethical considerations throughout the AI development lifecycle, developers can help prevent the misuse of AI technologies in election interference and other contexts (Bellamy et al., 2019).
@@
Developing robust legal and policy frameworks is essential to address the challenges posed by GenAI. These frameworks should encompass data protection laws, cybersecurity measures, and regulations specific to AI technologies. For instance, the General Data Protection Regulation (GDPR) in the European Union provides a strong foundation for data privacy and security, which is crucial in preventing unauthorized access and manipulation of data used in AI systems (Goodman and Flaxman, 2017).
Policies should also address the accountability of AI developers and users, ensuring that there are consequences for the misuse of AI technologies. This can include penalties for creating and disseminating deepfakes, as well as requirements for transparency in AI-generated content. Legal frameworks need to be adaptive and responsive to the rapidly evolving AI landscape, ensuring they remain effective in mitigating emerging threats (Mittelstadt et al., 2016).
In conclusion, addressing the societal implications of GenAI in election interference requires a comprehensive approach that integrates regulatory measures, technological solutions, public education, collaborative efforts, ethical development practices, and robust legal frameworks. By implementing these strategies, we can mitigate the risks associated with GenAI and protect the integrity of democratic processes (Ferrara, 2024b).

### v5

@@ Introduction
The constant exposure to AI-generated misinformation and deepfakes can have a psychological toll on the populace. The bombardment of conflicting information can lead to information overload, where individuals struggle to process and make sense of the vast amounts of data they encounter daily: This situation can cause increased stress, anxiety, and a sense of helplessness among individuals (fu2020social, ; matthes2020too, ; Brooks2024, ). Additionally, the fear of being deceived by AI-generated content can lead to a general mistrust of digital information. This mistrust can hinder the adoption of beneficial technologies and stifle innovation, as people become wary of engaging with digital platforms (ferrara2024genai, ).
Prevalence, Moral Panic, and Regulatory Targets
While concern over GenAI threats has grown significantly, the actual number of confirmed cases of election interference remains limited. This discrepancy raises the question of whether current discourse may be partially shaped by moral panic—the tendency to overestimate the immediate risks of novel technologies.
That said, GenAI is uniquely positioned to amplify threats at scale and speed, which justifies careful preemptive policy discussions. Vulnerable populations—such as older adults, individuals with low media literacy, or politically disengaged users—are more susceptible to synthetic content (ye2024susceptibility, ; luceri2025susceptibility, ). Their susceptibility makes them not only targets but potential amplifiers of GenAI misinformation.
From a policy perspective, regulatory frameworks should consider not only content creators and platforms but also the pathways of dissemination and the psychological traits that influence gullibility. Educational initiatives focused on digital literacy, platform accountability, and cross-sector collaboration will be vital in building long-term societal resilience.
Applications in Authoritarian Contexts
While the majority of documented GenAI election interference campaigns are observed in democratic or quasi-democratic contexts, the potential for exploitation in authoritarian regimes is substantial. However, empirical studies focused on authoritarian settings remain sparse—likely due to challenges in access, censorship, and the lack of media transparency.
Authoritarian governments may use GenAI not only to disrupt external democratic processes but also to maintain internal control. Tactics may include generating synthetic dissent to justify crackdowns, deploying AI-generated propaganda to simulate mass support, or discrediting foreign critics through deepfakes and forged documents.
The opacity of such regimes makes it difficult to quantify the scope and prevalence of GenAI use. This gap in the literature is a critical limitation, and future research must find creative and safe methodologies to illuminate the uses and abuses of AI technologies in these environments.
Table 3. Typology of GenAI Applications in Democratic Contexts
Modality Target Objective Example Use Case
Text (LLMs) Voters Disinformation Mass-produced false narratives on social media (weidinger2021ethical, )
Text (LLMs) Institutions Disruption AI-generated official-looking announcements or ballots (chiacchiaro2025generative, )
Audio Voters Discrediting Spoofed robocalls with fabricated candidate statements (schick2023toolformer, )
Video (Deepfakes) Candidates Discrediting Fake videos of misconduct during election campaigns (seymour2023beyond, ; marquart2023audiovisual, )
Memes (Image+Text) Voters Polarization Viral memes spreading divisive misinformation (pinto2024tracking, ; chen2025prevalence, )
Chatbots Voters Manipulation Impersonating campaign or electoral info services (li2024political, ; chiacchiaro2025generative, )
Table 4. Typology of GenAI Applications in Authoritarian Contexts
Modality Target Objective Example Use Case
Image (GANs) Voters Amplification Synthetic photos used in pro-regime propaganda (farrell2022spirals, ; gilding2023risking, )
Video (Deepfakes) Citizens Disruption Fabricated dissent to justify crackdowns (hameleers2023deepfake, ; li2022robust, )
Text (LLMs) Institutions Disinformation Ghostwritten op-eds simulating grassroots support (mcguffie2020radicalization, )
Interactive Media Voters Indoctrination Game-like simulations promoting regime narratives (birhane2023distort, )
@@
Effective regulation and oversight are crucial in governing the use of GenAI, particularly in sensitive areas like elections. Governments and international bodies need to establish clear guidelines and standards for the ethical use of AI technologies. These regulations should include provisions for transparency, accountability, and the prevention of misuse (floridi2019establishing, ).
One approach is the implementation of AI ethics guidelines, such as the European Union’s Ethics Guidelines for Trustworthy AI, which emphasize principles like human agency, technical robustness, privacy, and accountability (EUAIethics, ). Additionally, regulatory frameworks should mandate the disclosure of AI-generated content to ensure that the public can identify and scrutinize such content appropriately (openai2019release, ).
@@
Advances in technology can provide robust solutions to detect and mitigate the impact of AI-generated misinformation. For instance, digital watermarking and forensic techniques can be employed to identify deepfakes and other synthetic media (perov2020deepfacelab, ). These technologies can help verify the authenticity of digital content, making it easier to distinguish genuine from manipulated media.
Moreover, AI-driven detection systems can be developed to identify and neutralize botnets and fake accounts on social media platforms. These systems can use machine learning algorithms to detect patterns indicative of automated behavior, enabling platform operators to take swift action against malicious bots (yang2022botometer, ). Implementing such technologies requires collaboration between tech companies, researchers, and policymakers to ensure their effectiveness and scalability.
In response to the rising sophistication of GenAI, the research community has made progress in developing detection tools. For example, Salvi et al. (salvi2023timit, ) introduced the TIMIT-TTS dataset, a benchmark for evaluating text-to-speech detection methods, which helps advance audio-based synthetic media detection. Similarly, Hosler et al. (hosler2021deepfakes, ) proposed a semantic approach to deepfake detection that leverages emotional inconsistencies, offering a promising direction for identifying AI-generated visual content. These works underscore the importance of aligning detection technologies with the specific modality of GenAI content. As generative systems continue to evolve, multimodal detection techniques and adaptive classifiers will be crucial components of an effective defense strategy.
@@
Educating the public about the potential for AI-driven misinformation and how to critically evaluate information sources is crucial in combating the spread of false information. Public awareness campaigns can help individuals develop the skills needed to identify deepfakes, recognize signs of bot activity, and verify the credibility of news sources (rudin2019stop, ).
Educational initiatives should be integrated into school curricula and adult education programs to promote media literacy. These programs can teach critical thinking skills and provide tools for assessing the reliability of information online (ferrara2023social, ). Additionally, public service announcements and partnerships with media organizations can amplify these efforts, reaching a broader audience.
@@
Addressing the challenges posed by GenAI requires collaborative efforts among policymakers, technologists, and civil society. Multi-stakeholder initiatives can foster the development of best practices and shared standards for the ethical use of AI. For example, the Partnership on AI brings together diverse organizations to advance public understanding of AI and promote the responsible use of AI technologies (heer2018partnership, ).
Moreover, international cooperation is essential to address the global nature of online election interference. Governments and international organizations should work together to share information, coordinate responses, and develop joint strategies to counteract AI-driven threats. This collaboration can help build a unified front against the misuse of GenAI, ensuring that efforts are aligned and resources are effectively utilized (taddeo2018regulate, ).
@@
Promoting ethical AI development involves ensuring that AI systems are designed with fairness, accountability, and transparency in mind. Developers and researchers should adhere to ethical guidelines and consider the societal impact of their work. This includes conducting thorough risk assessments, implementing bias mitigation strategies, and fostering inclusive design practices (weidinger2022taxonomy, ).
Ethical AI development also involves creating AI systems that are interpretable and explainable. Interpretable models allow users to understand how decisions are made, increasing trust in AI systems and enabling more effective oversight (rudin2019stop, ). By prioritizing ethical considerations throughout the AI development lifecycle, developers can help prevent the misuse of AI technologies in election interference and other contexts (bellamy2019ai, ).
@@
Developing robust legal and policy frameworks is essential to address the challenges posed by GenAI. These frameworks should encompass data protection laws, cybersecurity measures, and regulations specific to AI technologies. For instance, the General Data Protection Regulation (GDPR) in the European Union provides a strong foundation for data privacy and security, which is crucial in preventing unauthorized access and manipulation of data used in AI systems (goodman2017european, ).
Policies should also address the accountability of AI developers and users, ensuring that there are consequences for the misuse of AI technologies. This can include penalties for creating and disseminating deepfakes, as well as requirements for transparency in AI-generated content. Legal frameworks need to be adaptive and responsive to the rapidly evolving AI landscape, ensuring they remain effective in mitigating emerging threats (mittelstadt2016ethics, ).
In conclusion, addressing the societal implications of GenAI in election interference requires a comprehensive approach that integrates regulatory measures, technological solutions, public education, collaborative efforts, ethical development practices, and robust legal frameworks. By implementing these strategies, we can mitigate the risks associated with GenAI and protect the integrity of democratic processes (ferrara2024risks, ).